OpenStudy (anonymous):

correct Sandra's problem. 2 + 3 = 2 + x. 2 2 + x = 2 - 3. x + 2 = -1. x = -3

OpenStudy (anonymous):

i can figure this out but i want to give a metal for whoever gets it first.

OpenStudy (anonymous):

its not hw or anything XD

Directrix (directrix):

I am having trouble reading the problem. Is that 22 corrrect or a typo? 2 + 3 = 2 + x. 2 2 + x = 2 - 3. x + 2 = -1. x = -3